    Author: Rodney Shaghoulian
    Github: github.com/RodneyShag
HackerRank: hackerrank.com/RodneyShag


Answer: 5/6

There are 6 possibilities on each die. On 2 dice, there are 6 * 6 = 36 possibilities

There are 6 cases where sum >= 10: (4,6), (5,5), (5,6), (6,4), (6,5), (6,6)

This gives us probability(sum >= 10) = 6/36 = 1/6

That means probability(sum <= 9) = 1 - 1/6 = 5/6
